OpenStudy (anonymous):

A $4,300.00 principal earns 5% interest, compounded semiannually. After 10 years, what is the balance in the account?

OpenStudy (anonymous):

it's all plug and chug.

OpenStudy (anonymous):

Well, would it be FV = PV (1+r)^n future value = 4300(1+0.05)^10

jimthompson5910 (jim_thompson5910):

close, that would be true if you were compounding annually

jimthompson5910 (jim_thompson5910):

instead you're compounding semiannually (so 2 times a year)

OpenStudy (anonymous):

So it would be 4300(1+0.05)^20

jimthompson5910 (jim_thompson5910):

closer

jimthompson5910 (jim_thompson5910):

you divide the 5% annual interest in half to get the interest rate per period (in this case, every 6 months) so 0.05/2 = 0.025

jimthompson5910 (jim_thompson5910):

it should be 4300(1+0.025)^20

jimthompson5910 (jim_thompson5910):

the 20 is the number of periods (2*10 = 20) basically 20 six-month periods

OpenStudy (anonymous):

7046.05?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

np